OpenStudy (anonymous):

Use the commutative property to rewrite the expression: 5 x + the quantity of negative 3 y

jimthompson5910 (jim_thompson5910):

the commutative property of addition allows you to add two numbers in any order

jimthompson5910 (jim_thompson5910):

for instance 1 + 2 is the same as 2 + 1

OpenStudy (anonymous):

Commutative property tells that a+b=b+a. so 5x+(3y)=3y+5x